# Opinion formation in Wikipedia Ising networks

### v2

@@ IV.1 Contest Trump, Putin, Xi Jinping
In Appendix Figures A2A3 we show the density, or frequency, of articles in the planes of color values $\eta_{Trump},\eta_{Putin},\eta_{Jinping}$ for the EN, RU, ZH editions that allows to see in a better way the distribution of articles and their color polarizations. (more technical details in the captions of these figures). Note that for the case of a pure two group competition these type of figures would give straight lines on the antidiagonal (from (0,1)(0,1)( 0 , 1 ) to (1,0)(1,0)( 1 , 0 )), since e.g. $1=\eta_{Trump}+\eta_{Putin}$ for a the pure Trump-Putin competition. In Appendix Figure A2, the data are indeed somewhat concentrated close to this antidiagonal showing the that modifications due to the influence of the third group of Jinping are rather modest for the editions of EN and RU.
@@ IV.3 Contest Liberalism, Communism, Nationalism
From the RGB map, we see that the strongest influence of Nationalism is for Serbia, Bosnia Herzegovina, Albania, Turkey, Romania, Bulgaria and in a less respect Pakistan, India, Bangladesh. The influence of Communism is rather local being represented only in Nepal. The influence of Liberalism is mostly spread over the world with the strongest adepts being UK, USA, Somaliland, Belgium, Netherlands and many others. The color of Russia is at the middle between Communism and Nationalism. This distribution of influence of the three concepts corresponds to the political orientations of these countries. Furthermore, in Appendix Figure A6 we present the histogram distribution in $\eta_{Liberalism}-\eta_{communism}$-plane. Here the articles are approximately distributed around the coordinate (0.33,0.23)(0.33,0.23)( 0.33 , 0.23 ) corresponding to $\eta_{Nationalism}=0.44$ which appears to be the typical polarization of these three concepts for 197 countries and also all other Wikipedia articles. In particular, here the distribution is not concentrated close to the antidiagonal as for certain other cases, showing that none of the three groups has a significantly reduced influence in comparison to the other two groups.
@@ V Effects of fluctuations at effective finite temperature
For the competition of two groups with different opinions (red vs, blue) the relation (1) for $Z_{i}$ determines the condition of spin updates with $\sigma_{i}=1$ if $Z_{i}>1$, $\sigma_{i}=-1$ if $Z_{i}-<0$ and no spin change if $Z_{i}=0$. Such a condition corresponds in the Monte Carlo process to the effective temperature $T=0$ since it gives a firm choice for the updated spin. It is interesting to analyze how stable this procedure is in presence of fluctuations produced by a finite effective temperature $T$. A finite $T$ value physically corresponds to the presence of finite probabilities $W_{\pm}(i)$ ($W_{+}(i)+W_{-}(i)=1$) to obtain the new spin value $\sigma_{i}=\pm 1$.
@@
To model this situation we write $Z_{i}=Z_{+}(i)-Z_{-}(i)$ as a difference of two positive quantities $Z_{\pm}(i)=\sum_{j\neq i}V_{ij}\delta_{\sigma_{j},\pm 1}\geq 0$ (i.e. sum only over all $j$ with either $\sigma_{j}=1$ or $\sigma=-1$ for the two cases $+$ or -- respectively). This is similar to the color score $Z_{i}(C)$ used in (2) if we use only two colors for spins ±1$\pm$ 1± 1. Then the probabilities $W_{\pm}(i)$ are determined by the relations
@@
The results for this finite temperature model of fluctuations are shown in Figures 12 and 13. Figure 12 shows that at $T\leq 0.75$ the density distribution of the (network) fraction of red nodes $f_{r}$ in at maximal iteration time $\tau=20$ is concentrated to the two regions $f_{r}\approx 0$ and $f_{r}\approx 1$ (similarly as Figure 1 which corresponds to $T=0$). In contrast at $T=1$ this density has a broad homogeneous distribution approximately in the range $0.2\leq f_{r}\leq 0.8$ and at $T=4$ all density is located in a narrow range around $f_{r}=0.5$. Indeed, at such a high temperature the probabilities to have spin up or down from (3) are very close $W_{+}\approx W_{-}\approx 0.5$ and hence we have approximately half od spins up and half down. The transition from a spin polarized steady-state at low temperatures to a non-polarized one takes place in the vicinity of a certain critical temperature $T_{c}$. Its value can be approximately determined by measuring the normalized number of spin flips or spin switches $N_{switch}$ at maximal iteration time $\tau=20$ as a function of temperature $T$. Up to now this quantity (at $T=0$) is essentially zero since at $\tau=20$ for a specific given pathway realization the spins of individual nodes are mostly in stable steady state.
@@ VI Discussion and conclusion
Appendix Table A1: Opinion polarization expressed by $\Delta\mu$ (following OPS for FR Wiki2024), for important personalities from French politics (top) and French richest persons (following Forbes top 10 ranking 2015-2024) (bottom). Emmanuel Macron corresponds to $\mu=1$ and Marine Le Pen corresponds to $\mu=-1$ with $\mu_{0}=-0.0252\pm 0.0023$ for $N_{r}=10^{5}$ and $\mu_{0}=-0.0279\pm 0.0008$ for $N_{r}=10^{6}$. The 2nd (3rd) column corresponds to data for $N_{r}=10^{5}$ ($N_{r}=10^{6}$) with a typical statistical error 0.0010.0010.001 (0.00030.00030.0003). The data for $N_{r}=10^{5}$ is statistically independent and not a subset of the data for $N_{r}=10^{6}$.

### v3

@@ IV.1 Contest Trump, Putin, Xi Jinping
In Appendix Figures A2A3 we show the density, or frequency, of articles in the planes of color values $\eta_{Trump},\eta_{Putin},\eta_{Jinping}$ for the EN, RU, ZH editions that allows to see in a better way the distribution of articles and their color polarizations. (more technical details in the captions of these figures). Note that for the case of a pure two group competition these type of figures would give straight lines on the antidiagonal (from $(0,1)$ to $(1,0)$), since e.g. $1=\eta_{Trump}+\eta_{Putin}$ for a the pure Trump-Putin competition. In Appendix Figure A2, the data are indeed somewhat concentrated close to this antidiagonal showing the that modifications due to the influence of the third group of Jinping are rather modest for the editions of EN and RU.
@@ IV.3 Contest Liberalism, Communism, Nationalism
From the RGB map, we see that the strongest influence of Nationalism is for Serbia, Bosnia Herzegovina, Albania, Turkey, Romania, Bulgaria and in a less respect Pakistan, India, Bangladesh. The influence of Communism is rather local being represented only in Nepal. The influence of Liberalism is mostly spread over the world with the strongest adepts being UK, USA, Somaliland, Belgium, Netherlands and many others. The color of Russia is at the middle between Communism and Nationalism. This distribution of influence of the three concepts corresponds to the political orientations of these countries. Furthermore, in Appendix Figure A6 we present the histogram distribution in $\eta_{Liberalism}-\eta_{communism}$-plane. Here the articles are approximately distributed around the coordinate $(0.33,0.23)$ corresponding to $\eta_{Nationalism}=0.44$ which appears to be the typical polarization of these three concepts for 197 countries and also all other Wikipedia articles. In particular, here the distribution is not concentrated close to the antidiagonal as for certain other cases, showing that none of the three groups has a significantly reduced influence in comparison to the other two groups.
@@ V Effects of fluctuations at effective finite temperature
For the competition of two groups with different opinions (red vs, blue) the relation (1) for $Z_{i}$ determines the condition of spin updates with $\sigma_{i}=1$ if $Z_{i}>0$, $\sigma_{i}=-1$ if $Z_{i}<0$ and no spin change if $Z_{i}=0$. Such a condition corresponds in the Monte Carlo process to the effective temperature $T=0$ since it gives a firm choice for the updated spin. It is interesting to analyze how stable this procedure is in presence of fluctuations produced by a finite effective temperature $T$. A finite $T$ value physically corresponds to the presence of finite probabilities $W_{\pm}(i)$ ($W_{+}(i)+W_{-}(i)=1$) to obtain the new spin value $\sigma_{i}=\pm 1$.
@@
To model this situation we write $Z_{i}=Z_{+}(i)-Z_{-}(i)$ as a difference of two positive quantities $Z_{\pm}(i)=\sum_{j\neq i,\,\sigma_{j}=\pm 1}V_{ij}\geq 0$ (i.e. sum only over all $j$ with either $\sigma_{j}=1$ or $\sigma_{j}=-1$ for the two cases ++ or - respectively). This is similar to the color score $Z_{i}(C)$ used in (2) if we use only two colors for spins $\pm 1$. Then the probabilities $W_{\pm}(i)$ are determined by the relations
@@
The results for this finite temperature model of fluctuations are shown in Figures 12 and 13. Figure 12 shows that at $T\leq 0.75$ the density distribution of the (network) fraction of red nodes $f_{r}$ in at maximal iteration time $\tau=20$ is concentrated to the two regions $f_{r}\approx 0$ and $f_{r}\approx 1$ (similarly as Figure 1 which corresponds to $T=0$). In contrast at $T=1$ this density has a broad homogeneous distribution approximately in the range $0.2\leq f_{r}\leq 0.8$ and at $T=4$ all density is located in a narrow range around $f_{r}=0.5$. Indeed, at such a high temperature the probabilities to have spin up or down from (3) are very close $W_{+}\approx W_{-}\approx 0.5$ and hence we have approximately half of spins up and half down. The transition from a spin polarized steady-state at low temperatures to a non-polarized one takes place in the vicinity of a certain critical temperature $T_{c}$. Its value can be approximately determined by measuring the normalized number of spin flips or spin switches $N_{switch}$ at maximal iteration time $\tau=20$ as a function of temperature $T$. Up to now this quantity (at $T=0$) is essentially zero since at $\tau=20$ for a specific given pathway realization the spins of individual nodes are mostly in stable steady state.
@@ VI Discussion and conclusion
Appendix Table A1: Opinion polarization expressed by $\Delta\mu$ (following OPS for FR Wiki2024), for important personalities from French politics (top) and French richest persons (following Forbes top 10 ranking 2015-2024) (bottom). Emmanuel Macron corresponds to $\mu=1$ and Marine Le Pen corresponds to $\mu=-1$ with $\mu_{0}=-0.0252\pm 0.0023$ for $N_{r}=10^{5}$ and $\mu_{0}=-0.0279\pm 0.0008$ for $N_{r}=10^{6}$. The 2nd (3rd) column corresponds to data for $N_{r}=10^{5}$ ($N_{r}=10^{6}$) with a typical statistical error $0.001$ ($0.0003$). The data for $N_{r}=10^{5}$ is statistically independent and not a subset of the data for $N_{r}=10^{6}$.